# Cycle-Consistent Bridge Diffusion Model for Accelerated MRI Reconstruction

### v1

@@ Abstract
Accelerated MRI reconstruction techniques aim to reduce examination time while maintaining high image fidelity, which is highly desirable in clinical settings for improving patient comfort and hospital efficiency. Existing deep learning methods typically reconstruct images from under-sampled data with traditional reconstruction approaches, but they still struggle to provide high-fidelity results. Diffusion models show great potential to improve fidelity of generated images in recent years. However, their inference process starting with a random Gaussian noise introduces instability into the results and usually requires thousands of sampling steps, resulting in sub-optimal reconstruction quality and low efficiency. To address these challenges, we propose Cycle-Consistent Bridge Diffusion Model (CBDM). CBDM employs two bridge diffusion models to construct a cycle-consistent diffusion process with a consistency loss, enhancing the fine-grained details of reconstructed images and reducing the number of diffusion steps. Moreover, CBDM incorporates a Contourlet Decomposition Embedding Module (CDEM) which captures multi-scale structural texture knowledge in images through frequency domain decomposition pyramids and directional filter banks to improve structural fidelity. Extensive experiments demonstrate the superiority of our model by higher reconstruction quality and fewer training iterations, achieving a new state of the art for accelerated MRI reconstruction in both fastMRI and IXI datasets.
@@
Accelerated MRI reconstructionn , Cycle-Consistent Bridge Diffusion , Contourlet Decomposition Embedding
@@ 1 Introduction
Magnetic resonance imaging (MRI) is widely used in clinical practice since it can highlight soft tissues with diverse contrasts. However, MRI scanning is slow ((e.g.,) a 1.5T brain examination typically takes 35-45 minutes [32]), leading to high costs and more motion artifacts [42]. Therefore, accelerated MRI reconstruction attracts great attention and various deep learning-based approaches have been proposed to reduce the scanning time and improve the reconstruction quality [20, 22, 11].
In general, accelerated MRI reconstruction aims to exploit fast-scanned MRI sequences, that are under-sampled in the k-space, and then transform them to full-sampled high-quality MRI data. Given the physical foundation in MRI reconstruction, data-driven works are often incorporated with traditional reconstruction operations to enhance both the model interpretability and imaging process. For example, Aggarwal et al. [1], Sun et al. [37], Zhang and Ghanem [43], Cui et al. [9], Hou and Li [17] suggest converting the iterative optimization process used in traditional compressed sensing (CS) techniques into trainable neural network. Similarly, Sriram et al. [36], Lv et al. [27], Knoll et al. [24] integrate traditional parallel imaging methods into deep neural networks to improve performance. Generative Adversarial Networks (GANs) [12] are also proposed to facilitate MRI reconstruction, boosting the generation performance [19, 27, 41, 8]. For instance, Chen et al. [6] enhanced the discriminative ability of GANs using wavelet transforms to improve MRI reconstruction accuracy. Cole et al. [8] implemented an unsupervised MRI reconstruction using GANs. Diffusion models are further used to accelerate MRI reconstruction, improving image fidelity. Xie and Li [39], Cao et al. [5], Hou et al. [18] developed diffusion models incorporating measurement and high-frequency priors, reducing unnecessary diffusion regions to enhance the generation quality.
Existing methods typically require multi-coil complex-valued data for accelerated MRI reconstruction, which is intermediate data processed within MRI scanners and thus is impractical to obtain in practice. In contrast, MRI magnitude images can be easily collected [38], and magnitude-based deep learning reconstruction can further significantly reduce MRI scanning time [40]. Therefore, we focus on the task of MRI accelerated reconstruction based on magnitude images, aiming to transform blurry and artifact-ridden MRI images into clear and artifact-free ones. As for accelerated MRI reconstruction based on magnitude images , GAN-based works such as pix2pix [21], cGAN [48] and pGAN [10] have achieved great progress. Then, diffusion-based methods [35, 15] further outperform GANs and improve the reconstruction quality. However, traditional diffusion models (e.g., DDPM [15] and SDE [35]) face significant challenges in clinical deployments. Their inference process, which begins with random Gaussian noise, often leads to result instability and requires thousands of sampling steps, resulting in suboptimal reconstruction quality and low efficiency.
Figure 2: The framework of the CBDM involves cycle-consistent diffusion between deterministically under-sampled and fully-sampled MR magnitude images. (a) A bridge diffusion model. (b) Forward cyclical consistency loss. (c) Backward cyclical consistency loss.
To address the above challenge, we first introduce a bridge diffusion model [25, 47], where the diffusion process is conducted between two deterministic images and has a clear guarantee to yield the desired conditional distribution without solely relying on a conditional image. To further enhance fine-grained details in reconstructed images and reduce diffusion time steps, we propose a cycle-consistent bridge diffusion method combined with a cycle-consistency loss. This method effectively improves the reconstruction quality of MRI images while significantly reducing the number of diffusion steps. On the other hand, the main difference between under-sampled MR magnitude images and fully-sampled MR magnitude images lies in the structural blurring and artifacts caused by the absence of high-frequency components. Although the iterative sampling process in diffusion models can implicitly capture structural information to a certain extent, it is still insufficient. Therefore, to explicitly account for the structural texture knowledge of MRI images, a Contourlet Decomposition Embedding Module (CDEM) is introduced. By utilizing iterative Laplacian pyramids [4] and directional filter banks [2, 7] to decompose the underlying features, this module extracts structural texture knowledge from each time-step image and integrates it into the denoising network, hence enhancing the network’s ability to capture structural textures.
Overall, the contributions of this paper are three-fold:
@@
    We propose a Cycle-Consistent Bridge Diffusion Model (CBDM), which formulates accelerated MRI reconstruction as a cycle-consistent diffusion process between under-sampled and fully-sampled MRI images. This process incorporates a cycle-consistency loss to preserve fine-grained details and reduce the number of diffusion time steps.
@@
    We propose a contourlet decomposition embedding module that leverages structural texture knowledge extracted from the frequency domain to mitigate the structural discrepancies caused by high-frequency misalignments between under-sampled and fully-sampled MR images.
@@
    Extensive experiments on the public fastMRI and IXI datasets demonstrate the effectiveness of our proposed Cycle-Consistent Bridge Diffusion-based Model for in-distribution and out-of-distribution accelerated MRI reconstruction.
2 Related works
@@ 2.1 Accelerated MRI Reconstruction
With the success of data-driven approaches, deep learning has been widely utilized in accelerated MRI reconstruction. Lv et al. [27], Sriram et al. [36] integrate deep learning with parallel imaging algorithms. [36] combined deep learning with the GRAPPA reconstruction algorithm. [44] transform the Iterative Shrinkage-Thresholding Algorithm (ISTA) [3] into a deep neural network to optimize compressive sensing (CS) for accelerated MRI reconstruction. [33] utilized an adversarial neural network to generate the missing k-space data. The generated and missing k-space data are subsequently transformed via Fourier transform and then fed into the discriminator for evaluation. [8] only requires under-sampled k-space data. The generator produces under-sampled k-space data from the generated image and the discriminator evaluates the k-space measurements instead of the MR image and provides the learned gradients back to the generator. Due to the high-fidelity generation characteristics of diffusion models, they have been widely applied in accelerated MRI reconstruction in recent years. [29] introduced an unconditional diffusion model, which is trained to generate coil-combined MR image samples derived from fully-sampled data. [14] proposed a conditional diffusion model for accelerated MRI reconstruction based on adaptive diffusion priors. To fully utilize the invariant nature of the low-frequency region in k-space during accelerated sampling, [5] introduced a conditional diffusion model based on high-frequency diffusion. Similarly, [39] proposed a measurement-conditioned diffusion model, which avoids diffusion in the undersampled mask region rather than only outside the low-frequency area, thereby preserving more prior information.
Although diffusion models in accelerated MRI reconstruction have evolved from unconditional to conditional and further to prior-based conditional diffusion, they typically set the ending states of the forward diffusion process to pure or partial Gaussian noise instead of a deterministic under-sampled MR image. Differently, our approach leverages the bridge diffusion model to achieve a deterministic image-to-image translation. Then, we propose the cycle-consistent constraint for the model training, further enhancing accelerated MRI reconstruction.
@@ 2.2 Diffusion Model
Recent advancements in diffusion models, such as those by [15, 28, 35], have significantly enhanced image generation, surpassing the performance of GANs [12]. These successes are attributed to key design choices, including network architecture [35, 23], optimized noise schedules [28, 23], improved sampling techniques [34, 26, 45], and advanced guidance methods [28, 16].
Bridge diffusion models [25, 47] have been proposed for image translation tasks, where the diffusion process is constructed between paired images. This approach starts reverse diffusion from a deterministic image (without noise), making the sampling process more stable. Inspired by the bridge diffusion model, we construct a cycle-consistent bridge diffusion model between under-sampled and fully-sampled MR magnitude images.
3 Method
Fig. 2 shows an overview of our Cycle-Consistent Bridge Diffusion Model, including two bridge diffusion models for transformations $\boldsymbol{x}\leftarrow\boldsymbol{y}$ and $\boldsymbol{y}\leftarrow\boldsymbol{x}$, respectively. Here, $\boldsymbol{x}_{0}$ and $\boldsymbol{y}_{0}$ denote the fully-sampled and under-sampled magnitude images. $\to$ and $\leftarrow$ denote the forward and reverse processes in each diffusion model, respectively.
@@ 3.1 Bridge Diffusion Model
3.2 Cycle-Consistent Bridge Diffusion Model
To enhance the diffusion model training, we further adopt a cyclical design. Specifically, Equation (6) indicates that, during training, the target image $\boldsymbol{x}_{0}$ or $\boldsymbol{y}_{0}$ can be estimated by a noisy state $\boldsymbol{x}_{t1}$ and $\boldsymbol{y}_{t1}$, respectively. Therefore, the reconstructed fully-sampled and under-sampled MR magnitude images $\boldsymbol{\bar{x}}_{0}$ and $\boldsymbol{\bar{y}}_{0}$ become:
@@ 3.2 Cycle-Consistent Bridge Diffusion Model
We then set the reconstructed ones as new end states. Hence, the intermediate states $\boldsymbol{x}^{\prime}_{t2}$ and $\boldsymbol{y}^{\prime}_{t2}$ become:
@@
Similarly, the reverse processes $\boldsymbol{y}\leftarrow\boldsymbol{\bar{x}}_{0}$ and $\boldsymbol{x}\leftarrow\boldsymbol{\bar{y}}_{0}$ can obtain the new reconstructed images $\boldsymbol{\bar{\bar{y}}}_{0}$ and $\boldsymbol{\bar{\bar{x}}}_{0}$, respectively as:
@@
Therefore, a cycle-consistency loss can be used to reduce the distance between $\boldsymbol{x}_{0}$ and $\boldsymbol{\bar{\bar{x}}}_{0}$, and between $\boldsymbol{y}_{0}$ and $\boldsymbol{\bar{\bar{y}}}_{0}$, respectively as:
$\mathcal{L}_{\boldsymbol{x}}^{con}(\theta 1,\theta 2)=\mathbb{E}_{\boldsymbol{x}_{0},\boldsymbol{y}_{0}}[||\boldsymbol{x}_{0}-\boldsymbol{\bar{\bar{x}}}_{0}||]$ (10)
$\mathcal{L}_{\boldsymbol{y}}^{con}(\theta 1,\theta 2)=\mathbb{E}_{\boldsymbol{x}_{0},\boldsymbol{y}_{0}}[||\boldsymbol{y}_{0}-\boldsymbol{\bar{\bar{y}}}_{0}||].$
Overall, the total loss is a weighted sum of construction loss and consistency losses as
$\mathcal{L}_{total}=\lambda(\mathcal{L}_{\boldsymbol{x}}^{con}+\mathcal{L}_{\boldsymbol{y}}^{con})+(\mathcal{L}_{\boldsymbol{x}\leftarrow\boldsymbol{y}}^{rec}+\mathcal{L}_{\boldsymbol{y}\leftarrow\boldsymbol{x}}^{rec})$ (11)
Algorithm 1 Pseudo training codes of our cycle-consistent bridge diffusion model for accelerated MRI reconstruction
@@
12:     Forward diffusion $\boldsymbol{y}_{0}\to\boldsymbol{\bar{x}_{0}}$
@@
15:     Forward diffusion $\boldsymbol{x}_{0}\to\boldsymbol{\bar{y}_{0}}$

### v2

# Self-Consistent Nested Diffusion Bridge for Accelerated MRI Reconstruction

@@ Abstract
Accelerated MRI reconstruction plays a vital role in reducing scan time while preserving image quality. While most existing methods rely on complex-valued image-space or k-space data, these formats are often inaccessible in clinical practice due to proprietary reconstruction pipelines, leaving only magnitude images stored in DICOM files. To address this gap, we focus on the underexplored task of magnitude-image-based MRI reconstruction. Recent advancements in diffusion models, particularly denoising diffusion probabilistic models (DDPMs), have demonstrated strong capabilities in modeling image priors. However, their task-agnostic denoising nature limits performance in source-to-target image translation tasks, such as MRI reconstruction. In this work, we propose a novel Self-Consistent Nested Diffusion Bridge (SC-NDB) framework that models accelerated MRI reconstruction as a bi-directional image translation process between under-sampled and fully-sampled magnitude MRI images. SC-NDB introduces a nested diffusion architecture with a self-consistency constraint and reverse bridge diffusion pathways to improve intermediate prediction fidelity and better capture the explicit priors of source images. Furthermore, we incorporate a Contour Decomposition Embedding Module (CDEM) to inject structural and textural knowledge by leveraging Laplacian pyramids and directional filter banks. Extensive experiments on the fastMRI and IXI datasets demonstrate that our method achieves state-of-the-art performance compared to both magnitude-based and non-magnitude-based diffusion models, confirming the effectiveness and clinical relevance of SC-NDB.
@@
Accelerated MRI Reconstruction , Nested Diffusion Bridge , Self-Consistent , Contour Decomposition
journal: Pattern Recognition
@@ 1 Introduction
Magnetic resonance imaging (MRI) is widely used in clinical practice due to its ability to visualize various human regions with diverse contrasts. However, MRI scanning is inherently slow (e.g., a 3.0T brain examination typically takes 35–45 minutes [34]), resulting in high costs and increased susceptibility to motion artifacts [42]. Consequently, accelerated MRI reconstruction has garnered significant attention, leading to the development of various deep learning-based approaches aimed at reducing scanning time and improving reconstruction quality [22, 23, 12, 38, 7, 31, 28].
Accelerated MRI reconstruction methods can be categorized based on the type of input data into image-space, k-space, and magnitude-image approaches. The distinction between magnitude images and raw data is illustrated in Fig. 3. While image-space and k-space data can be mutually converted through Fourier transform, magnitude-image cannot be reversibly transformed back to either image-space or k-space data.
Most accelerated MRI reconstruction methods rely on complex-valued data (k-space or image-space) [21, 7, 43, 13, 8, 20, 15, 16, 40, 6, 19, 46, 11], which is derived from intermediate processes during MRI scanning. However, obtaining such raw data is clinically impractical since manufacturer-specific reconstruction pipelines are inherently required before DICOM file storage, which only preserves magnitude images. In practical scenarios, deep learning-based MRI accelerated reconstruction using magnitude-image [39] is also widely adopted in clinical applications, offering better generalization and processing speed [41]. Therefore, this study primarily focuses on the relatively underexplored area of magnitude-image-based accelerated MRI reconstruction.
Many studies has adopted denoising diffusion probabilistic models (DDPMs) to capture explicit image priors [17, 30, 37, 36], enhancing training stability and improving image fidelity. In DDPMs, the forward diffusion process gradually degrades the target image by repeatedly adding Gaussian noise until it becomes pure noise. Starting from a randomly sampled noise image, the reverse diffusion process then progressively removes the noise using a neural network to recover the target image, conditioned on the source image. However, DDPMs learn a task-agnostic denoising transformation from noise to the target image, which can weaken the guidance from the source image and thus may underperform in source-to-target image translation tasks.
Recently, the bridge diffusion model has emerged as a promising approach for enhancing source image relevance in image translation tasks by enabling direct transformations between two distinct distribution images [25, 26, 47]. To this end, bridge diffusion defines the start and end points of the forward diffusion process based on the source and target images, respectively. Sampling begins directly from the source images, and the reverse process incrementally transforms it toward the target image. Unlike conventional bridge diffusion models,  Arslan et al. [1] introduces a novel noise scheduling strategy in the forward process, where the variance monotonically increases according to the end state corresponding to the noise-added source images. This design aims to capture soft priors from the source modality and improve generalization.
While existing bridge diffusion methods focus primarily on innovations in noise scheduling, limited attention has been paid to self-consistency in the reverse diffusion process during training. To address this, we propose a novel Self-Consistent Nested Diffusion Bridge (SC-NDB) to enhance source-to-target image translation performance. In contrast to conventional bridge diffusion models, SC-NDB incorporates both a reverse bridge diffusion model and a bi-directional translation process during training, aiming to improve the accuracy of intermediate state predictions. This allows the model to more accurately capture explicit priors from the source image, thus facilitating more effective information transfer between source and target images.
The difference between fully-sampled and under-sampled MRI images is primarily manifested in the high-frequency components. To explicitly incorporate structural and textural knowledge specific to MRI, we introduce a Contour Decomposition Embedding Module (CDEM) to further boost performance. This module decomposes low-level features from each time-step image using iterative Laplacian pyramids [5] and directional filter banks [2, 9], extracting structural and textural information and integrating it into the denoising network, thereby enhancing the network’s ability to preserve anatomical details.
Extensive validation was conducted on the publicly available fastMRI and IXI datasets. Our results clearly demonstrate the superior performance of SC-NDB compared to competitive magnitude-image-based diffusion models and non-magnitude-image-based models (k-space and image-space), confirming its effectiveness in MRI reconstruction tasks.
Overall, the contributions in this paper are three-fold:
@@
    We propose a Self-Consistent Nested Diffusion Bridge (SC-NDB), which formulates accelerated MRI reconstruction as a bi-directional translation process between under-sampled and fully-sampled MRI magnitude images. This process incorporates a self-consistency loss via a nested diffusion bridge to more accurately capture explicit priors from the source images.
@@
    We propose a Contourlet Decomposition Embedding Module (CDEM) that leverages structural texture knowledge extracted from the frequency domain to mitigate the structural discrepancies caused by high-frequency misalignment between under-sampled and fully-sampled MR magnitude images.
@@
    Extensive experiments on the public fastMRI and IXI datasets demonstrate the effectiveness of our proposed Self-Consistent Diffusion Bridge Model for in-distribution and out-of-distribution accelerated MRI reconstruction.
Figure 1: The framework of the SC-NDB involves two nested diffusion bridges between deterministically under-sampled and fully-sampled MR magnitude images.
2 Related Work
@@ 2.1 Accelerated MRI Reconstruction
With the success of data-driven approaches, deep learning has been widely utilized in accelerated MRI reconstruction. Lv et al. [29], Sriram et al. [38] integrate deep learning with parallel imaging algorithms. [38] combined deep learning with the GRAPPA reconstruction algorithm. [44] transform the Iterative Shrinkage-Thresholding Algorithm (ISTA) [4] into a deep neural network to optimize compressive sensing (CS) for accelerated MRI reconstruction. [35] utilized an adversarial neural network to generate the missing k-space data. The generated and missing k-space data are subsequently transformed via Fourier transform and then fed into the discriminator for evaluation. [10] only requires under-sampled k-space data. The generator produces under-sampled k-space data from the generated image and the discriminator evaluates the k-space measurements instead of the MR image and provides the learned gradients back to the generator. Due to the high-fidelity generation characteristics of diffusion models, they have been widely applied in accelerated MRI reconstruction in recent years. [31] introduced an unconditional diffusion model, which is trained to generate coil-combined MR image samples derived from fully-sampled data. [16] proposed a conditional diffusion model for accelerated MRI reconstruction based on adaptive diffusion priors. To fully utilize the invariant nature of the low-frequency region in k-space during accelerated sampling, [6] introduced a conditional diffusion model based on high-frequency diffusion. Similarly, [40] proposed a measurement-conditioned diffusion model, which avoids diffusion in the under-sampled mask region rather than only outside the low-frequency area, thereby preserving more prior information. To further enhance the guidance derived from the source images, bridge diffusion models have recently been introduced into the task of MRI accelerated reconstruction. Hu et al. [21] proposed a bridge diffusion model that adaptively calibrates an unknown forward model to reinforce measurement consistency throughout the sampling iterations, addressing the blind inverse problem in MRI accelerated reconstruction. Gao et al. [13] introduced a self-supervised bridge diffusion approach that directly trains on available noisy measurements without the need for any high-quality reference images.
Figure 3: Categorized reconstruction methods based on data types: image-space, k-space, and magnitude-image. Our proposed method specifically focuses on the magnitude image scenario.
Although diffusion models for accelerated MRI reconstruction have evolved from conditional diffusion to bridge diffusion conditioned on source images, they are typically built upon complex-valued image-space or k-space data as source and target domains. In contrast, diffusion modeling based on magnitude images (widely stored in clinical DICOM formats) has been rarely explored.
@@ 2.2 Diffusion Model
Recent advancements in diffusion models, such as those by [17, 30, 37, 3], have significantly enhanced image generation, surpassing the performance of GANs [14]. These successes are attributed to key design choices, including network architecture [37, 24], optimized noise schedules [30, 24], improved sampling techniques [36, 27, 45], and advanced guidance methods [30, 18].
Bridge diffusion models [25, 47, 26, 1] have been proposed for image translation tasks, where the diffusion process is constructed between paired images. This approach starts reverse diffusion from a deterministic image (without noise), making the sampling process more stable. Existing studies on diffusion bridges have primarily focused on innovations in noise scheduling, while overlooking the self-consistency of intermediate states. This limitation hinders the accurate capture of prior distributions from the source images. To address this, we propose a Self-Consistent Nested Diffusion Bridge (SC-NDB) to model the translation from under-sampled MRI images to fully-sampled MRI images.
3 Methodology
Fig. 1 provides an overview of our Self-Consistent Nested Diffusion Bridge, where two nested diffusion bridges, A and B, are constructed using four distinct diffusion bridges. Specifically, the diffusion bridges $\boldsymbol{x}\leftarrow\boldsymbol{y}$ and $\boldsymbol{x}\leftarrow\boldsymbol{\overline{y}}$ share parameters $\theta_{1}$, while $\boldsymbol{y}\leftarrow\boldsymbol{x}$ and $\boldsymbol{y}\leftarrow\boldsymbol{\overline{x}}$ share parameters $\theta_{2}$. Here, $\boldsymbol{x}_{0}$ and $\boldsymbol{y}_{0}$ denote the fully-sampled and under-sampled magnitude images. $\to$ and $\leftarrow$ denote the forward and reverse processes in each diffusion model, respectively.
@@ 3.1 Bridge Diffusion Model
3.2 Self-Consistent Nested Diffusion Bridge
To more accurately capture the prior distribution of source images, a new diffusion bridge is nested to enforce self-consistency. Specifically, Equation (6) indicates that, during training, the target image $\boldsymbol{x}_{0}$ or $\boldsymbol{y}_{0}$ can be estimated by a noisy state $\boldsymbol{x}_{t1}$ and $\boldsymbol{y}_{t1}$, respectively. Therefore, the reconstructed fully-sampled and under-sampled MR magnitude images $\boldsymbol{\bar{x}}_{0}$ and $\boldsymbol{\bar{y}}_{0}$ become:
@@ 3.2 Self-Consistent Nested Diffusion Bridge
We then designate the reconstructed samples as the new end states of the new diffusion bridges $\boldsymbol{y}\leftarrow\boldsymbol{\overline{x}}$ and $\boldsymbol{x}\leftarrow\boldsymbol{\overline{y}}$, respectively. Consequently, the intermediate states $\boldsymbol{x}^{\prime}_{t2}$ and $\boldsymbol{y}^{\prime}_{t2}$ are defined as follows:
@@
Similarly, the reverse processes of the new diffusion bridges $\boldsymbol{y}\leftarrow\boldsymbol{\bar{x}}$ and $\boldsymbol{x}\leftarrow\boldsymbol{\bar{y}}$ result in the newly reconstructed images $\boldsymbol{\bar{\bar{y}}}_{0}$ and $\boldsymbol{\bar{\bar{x}}}_{0}$, respectively. These are given by:
@@
Therefore, a self-consistent loss can be used to reduce the distance between $\boldsymbol{x}_{0}$ and $\boldsymbol{\bar{\bar{x}}}_{0}$, and between $\boldsymbol{y}_{0}$ and $\boldsymbol{\bar{\bar{y}}}_{0}$, respectively as:
$\mathcal{L}_{\boldsymbol{x}}^{selfcon}(\theta 1,\theta 2)=\mathbb{E}_{\boldsymbol{x}_{0},\boldsymbol{y}_{0}}[||\boldsymbol{x}_{0}-\boldsymbol{\bar{\bar{x}}}_{0}||]$ (10)
$\mathcal{L}_{\boldsymbol{y}}^{selfcon}(\theta 1,\theta 2)=\mathbb{E}_{\boldsymbol{x}_{0},\boldsymbol{y}_{0}}[||\boldsymbol{y}_{0}-\boldsymbol{\bar{\bar{y}}}_{0}||].$
Algorithm 1 Pseudo training codes of our Self-Consistent Nested Diffusion Bridge for accelerated MRI reconstruction
@@
12:     New Forward diffusion $\boldsymbol{y}_{0}\to\boldsymbol{\bar{x}_{0}}$
@@
15:     New Forward diffusion $\boldsymbol{x}_{0}\to\boldsymbol{\bar{y}_{0}}$
@@
Figure 4: Visual comparison of sparsity representations using (a) traditional Wavelet Transform and (b) Contourlet Transform.
Overall, the total loss is a weighted sum of reconstruction loss and self-consistent loss as
$\mathcal{L}_{total}=\lambda(\mathcal{L}_{\boldsymbol{x}}^{selfcon}+\mathcal{L}_{\boldsymbol{y}}^{selfcon})+(\mathcal{L}_{\boldsymbol{x}\leftarrow\boldsymbol{y}}^{rec}+\mathcal{L}_{\boldsymbol{y}\leftarrow\boldsymbol{x}}^{rec})$ (11)